# Language Models That Walk the Talk: A Framework for Formal Fairness Certificates

### v1

@@ Appendix A Zonotope-based verification of neural networks
Given a center vector $c\in\mathbb{R}^{n}$ and a generator matrix $G\in\mathbb{R}^{n\times q}$, a zonotope girard2005reachability is defined as
@@
The input set $\mathcal{X}$ should capture all points $x$ for which $\psi_{\text{in}}(x)$ holds. Typically, $\mathcal{X}$ captures the $\ell_{\infty}$-bounded perturbations around some given input $x_{0}\in\mathbb{R}^{n}$. Given a perturbation radius ϵ+italic-$\epsilon\in\mathbb{R}_{+}$, the input zonotope is given by
@@ B.1 Gender-specific dataset
The gender-specific word list used for pre-training is sourced from bolukbasi2016man , where the authors derived gendered terms from WordNet based on definitional criteria. We adopt this word set and apply the Word Embedding Association Test (WEAT) to categorize terms into male- and female-associated groups used predefined attribute set (Table 4), this approach follows established practice, as is previously used by literature studying gender bias in embeddings or corpus betti2023large ; Caliskan_2017 ; charlesworth2021gender .
Table 4: List of attribute word sets used in the Word Embedding Association Test (WEAT), originally derived from Caliskan_2017 ; betti2023large and subsequently expanded by the authors to enrich semantic coverage and improve gender-related classification.
@@ B.2 Details on contrastive loss function
The loss function is designed to jointly optimize the embedding space by drawing semantically similar word pairs closer together while pushing dissimilar pairs apart. For similar pairs, a linear pull term proportional to the predicted distance is introduced, promoting tighter semantic clustering. To ensure numerical stability and prevent excessively large gradients, we deliberately avoid logarithmic or high-degree penalty functions, opting instead for a linear formulation. For dissimilar pairs, we employ a standard contrastive loss component, which penalizes instances where the predicted distance falls below a predefined margin, thereby enforcing a minimum separation between unrelated terms.555 While pre-training aims to minimize the $\ell_{2}$ distance between embeddings, the verification process instead focuses on finding the maximum verifiable radius under the $\ell_{\infty}$ norm, since formal verification over the $\ell_{2}$ domain is not directly possible using zonotopes. The authors intentionally adopt the $\ell_{2}$ distance for pre-training because it minimizes the overall distance between points, rather than just the maximum per-dimension difference. This encourages the model to learn more generalized and robust representations. Furthermore, the authors observed in their experiments that after pre-training, the $\ell_{2}$ distance between semantically similar words (such as gendered word pairs) is often slightly larger than their corresponding $\ell_{\infty}$ distance (as can be seen in Section C.3.4). This supports the idea that utilizing the $\ell_{2}$ norm benefits downstream verification..
@@ B.3 Embedding model pre-training data construction
Given these limitations and the impracticality of verifying large-scale contextualized models like BERT shi2020robustness ; bonaert2021fast , we pre-train a Keras embedding layer ketkar2017introduction . To construct an effective training set for contrastive loss optimization, we curated word pairs based on the semantic gender relationship using a heuristic approach. Each male-labeled word was explicitly paired with the terms woman and female to anchor the gender axis, and further matched with a diverse sample of 50-100 female-labeled words to form similar (label=0) pairs. This ensures that embeddings of gender-related terms are pulled closer in the vector space. To balance the contrastive objective, we generated dissimilar (label=1) pairs by selecting words from a broad English vocabulary that have no direct lexical or semantic relationship (e.g., synonyms, derivations) to any gender-specific term. This distinction was enforced by filtering out related terms using WordNet relations such as synonyms, hypernyms, and derivationally related forms.
@@
To enhance the general semantic capacity of the embedding space, particularly because the embedding layer is later frozen during task-specific backpropagation, we conducted an auxiliary contrastive training phase focused on common semantic relationships such as synonymy [^6] This general semantic training step proved essential, as model performance degraded significantly in its absence. This outcome is expected, given that the embedding layer, when integrated within the transformer-based architecture, is frozen during downstream training and therefore does not receive gradient updates to refine its semantic representations.. Using WordNet, we generated word pairs labeled as similar (label=0) by extracting synonym pairs, while carefully excluding any gender-specific terms to preserve the integrity of gender representations learned in earlier pre-training stages, thereby maintaining the intended semantic distances between gendered terms. To form the dissimilar (label=1) pairs, we randomly sampled words from the general vocabulary that were not semantically related [^7] Each non-gendered word is paired with all valid WordNet synonyms, excluding any gendered terms, resulting in a variable number of synonym pairs per word. To ensure a balanced dataset, an equal number of dissimilar word pairs is generated, maintaining a one-to-one ratio with the total number of synonym pairs., again ensuring that no gender-specific tokens were involved. This allowed the model to learn to cluster semantically similar words and push apart unrelated ones. By training the embedding space on these general lexical relationships, we improve its ability to capture broader semantic structure, which is especially important since the embeddings are utilized as fixed features in downstream transformer-based language model text classification tasks. This step ensures that the embeddings are not only gender-aware but also semantically meaningful in a general linguistic context.
@@
For pre-training the toxicity-related embedding model, a similar pre-training strategy was employed—using a contrastive learning framework to pull semantically similar words closer and push dissimilar words apart [^8] To handle rare or profane words not explicitly found in the tokenizer vocabulary, we use BERT’s WordPiece-based subword tokenization. This strategy breaks out-of-vocabulary terms into smaller known units (e.g., mother-fucker becomes [mother, -, fuck, ##er]), allowing consistent representation throughout the embedding layer pre-training and language model training phases without full word-level coverage. These subwords are embedded and mean-pooled, enabling the model to capture semantic information even for obfuscated or stylized expressions.. The strength of the pulling and pushing mechanisms (i.e., the values of $\alpha$ and the margin), along with other hyperparameters, can be found in Section C.1. The distance between gendered words, toxicity words, synonyms, and unrelated words before and after pre-training of embedding can be found in Section C.3.4.
@@ B.4 Adult census dataset preprocessing
Transformers require one-dimensional token sequences as input, whereas tabular data is inherently two-dimensional badaro2023transformers ; however, the demographic information in the Adult Census dataset is structured in a tabular format, with each attribute represented as a separate column. To bridge this gap, the data and its contextual structure must be serialized before being processed by the model badaro2023transformers . A critical step in adapting tabular data for transformer-based models is converting the tabular structure into a linearized, one-dimensional format. Common strategies include: column-wise concatenation of values, insertion of special separator tokens between fields badaro2023transformers ; 10.1145/3514221.3517906 , and use of text templates to represent each row as a structured sentence chen2019tabfact ; suadaa2021towards ; chen2020logical . More recent approaches move beyond predefined templates by generating natural language sentences directly from tabular data. This is achieved by fine-tuning sequence-to-sequence models such as T5 raffel2020exploring or GPT radford2019language , as demonstrated in prior works thorne2021database ; neeraja2021incorporating . We adopt simple concatenation for its straightforward implementation and effectiveness in our setting.
To adapt the Adult Census dataset for use with a sentence-to-label transformer-based binary classification model, we converted the structured tabular data into natural language sentences. The original dataset contains both numerical and categorical features, along with information indicating whether an individual’s income exceeds \$50K. Since transformer models (e.g., BERT) operate on text sequences, we constructed a heuristic template that combines relevant features into coherent sentences. Specifically, attributes such as workclass, education, occupation, sex, marital status, relationship, race, native country, and hours per week were embedded into a natural language sentence format. The salary attribute, serving as the ground truth label for prediction, was intentionally excluded from the sentence construction. This transformation preserves the semantic content of the original features while enabling the use of sentence-based binary text classification models on structured tabular data. An example tabular data instance before and after sentence construction can be found in Table 5999The feature ”fnlwgt”, along with several other complex numerical attributes, was excluded from this experiment, as the authors deliberately chose to focus on leveraging textual, simple numerical, and categorical features for model training and representation..
@@ B.5 Data augmentation
Text classification is a fundamental task in natural language processing, and achieving high performance often depends heavily on the size and quality of the training data wei2019eda ; kobayashi2018contextual . Automatic data augmentation is widely employed in machine learning and computer vision, among other domains, to enhance model robustness, particularly when working with limited datasets 10.5555/645754.668381 ; szegedy2015going ; 10.1145/3065386 ; 7113823 ; ko2015audio .
A commonly used data augmentation strategy is synonym replacement, where selected words are substituted with their synonyms. This can be achieved using handcrafted lexical resources like WordNet 10.1145/219717.219748 ; zhang2016yelp , word similarity computations wang-yang-2015-thats , or contextual embedding-based methods kobayashi2018contextual .
We employed a simple WordNet-based synonym replacement strategy to augment the dataset in the salary prediction experiments, aiming to increase data volume and enhance classification accuracy. This approach led to a notable improvement in test accuracy, rising from 80% to around 90% [^10] We utilized straightforward WordNet synonym replacement, as empirical observations indicated that this method is already effective enough in improving model performance..
@@ C.1 Evaluation details
The selection of margin and $\alpha$ values is based on empirical observation. The goal is not only to ensure that gendered words are embedded in close proximity but also to promote the clustering of synonyms while pushing antonyms farther apart, without compromising the relative closeness of gendered word pairs. The margin values are also selected empirically based on the nuanced relationship between the embedding space and the resulting verifiable radius. A detailed study of this relationship is provided in Section C.3.2. The loss function used in this study is focal loss Lin_2017_ICCV , a training-time calibration technique that directly integrates class imbalance handling into the learning process. It works by adapting the loss function to emphasize challenging, misclassified samples.
To evaluate the fairness score of the models, we use the MATLAB toolbox CORA althoff2015introduction , which enables formal verification using set-based computing kochdumper2022open ; ladner2023automatic ; bonaert2021fast . Please visit Appendix A for a detailed explanation of this process using zonotopes.
@@
Model Salary Prediction & Gender Bias Mitigation
$\alpha$ Margin $\alpha$ (Gen. Sem. Train.) Margin (Gen. Sem. Train.) Epoch Epoch (Gen. Sem. Train.)
@@
Model Toxicity Detection & Content Moderation
$\alpha$ Margin $\alpha$ (Gen. Sem. Train.) Margin (Gen. Sem. Train.) Epoch Epoch (Gen. Sem. Train.)
@@
Model Salary Prediction & Gender Bias Mitigation
Vocab (BERT) Seq. Len. Embed Dim Heads FF Dim Dropout Epochs Batch
@@
Model Toxicity Detection & Content Moderation
Vocab Seq. Len. Embed Dim Heads FF Dim Dropout Epochs Batch
@@ C.2 Sentence-wise maximum verifiable perturbation radius
Model Salary Prediction & Gender Bias Mitigation (8 blocks)
Statistic Bounds
@@
Sentence index Max verifiable perturbed noise radii ($\uparrow$)
Noise Radii Sentence 1 9.6552e-04
Sentence 2 8.6406e-04
Sentence 3 9.2952e-04
Sentence 4 9.6552e-04
Sentence 5 9.5243e-04
Sentence 6 9.0333e-04
Sentence 7 9.5243e-04
Sentence 8 9.6552e-04
Sentence 9 9.2952e-04
Sentence 10 9.3115e-04
Sentence 11 8.5260e-04
Sentence 12 8.1660e-04
Sentence 13 9.6552e-04
Sentence 14 9.9661e-04
Sentence 15 7.4951e-04
Sentence 16 1.1161e-03
Sentence 17 4.9749e-04
Sentence 18 4.9749e-04
Sentence 19 3.9930e-04
Sentence 20 9.7043e-04
@@
Model Salary Prediction & Gender Bias Mitigation (8 blocks)
Statistic Bounds
Embedding Bounds Minimum ($\inf$) -1.696724
Maximum ($\sup$) 1.735214
Range ($\sup-\inf$) 3.431938
Sentence index Max verifiable perturbed noise radii ($\uparrow$)
Noise Radii Sentence 1 4.9393e-03
Sentence 2 7.0710e-03
Sentence 3 6.4731e-03
Sentence 4 1.1698e-03
Sentence 5 7.6429e-03
Sentence 6 6.1091e-03
Sentence 7 5.4592e-04
Sentence 8 7.3310e-03
Sentence 9 5.3552e-03
Sentence 10 6.4731e-03
Sentence 11 4.2894e-03
Sentence 12 8.4228e-03
Sentence 13 5.0433e-03
Sentence 14 5.2512e-03
Sentence 15 5.8492e-03
Sentence 16 5.3032e-03
Sentence 17 7.4089e-03
Sentence 18 2.5996e-03
Sentence 19 4.9393e-03
Sentence 20 4.0294e-03

### v2

@@ Appendix A Zonotope-based verification of neural networks
Given a center vector $c\in\mathbb{R}^{n}$ and a generator matrix $G\in\mathbb{R}^{n\times q}$, a zonotope (Girard, 2005) is defined as
@@
The input set $\mathcal{X}$ should capture all points $x$ for which $\psi_{\text{in}}(x)$ holds. Typically, $\mathcal{X}$ captures the $\ell_{\infty}$-bounded perturbations around some given input $x_{0}\in\mathbb{R}^{n}$. Given a perturbation radius $\epsilon\in\mathbb{R}_{+}$, the input zonotope is given by
@@ B.1 Gender-specific dataset
The gender-specific word list used for pre-training is sourced from (Bolukbasi et al., 2016), where the authors derived gendered terms from WordNet based on definitional criteria. We adopt this word set and apply the Word Embedding Association Test (WEAT) to categorize terms into male- and female-associated groups used predefined attribute set (Table 4), this approach follows established practice, as is previously used by literature studying gender bias in embeddings or corpus (Betti et al., 2023; Caliskan et al., 2017b; Charlesworth et al., 2021).
Table 4: List of attribute word sets used in the Word Embedding Association Test (WEAT), originally derived from (Caliskan et al., 2017b; Betti et al., 2023) and subsequently expanded by the authors to enrich semantic coverage and improve gender-related classification.
@@ B.2 Details on contrastive loss function
The loss function is designed to jointly optimize the embedding space by drawing semantically similar word pairs closer together while pushing dissimilar pairs apart. For similar pairs, a linear pull term proportional to the predicted distance is introduced, promoting tighter semantic clustering. To ensure numerical stability and prevent excessively large gradients, we deliberately avoid logarithmic or high-degree penalty functions, opting instead for a linear formulation. For dissimilar pairs, we employ a standard contrastive loss component, which penalizes instances where the predicted distance falls below a predefined margin, thereby enforcing a minimum separation between unrelated terms.
While pre-training aims to minimize the $\ell_{2}$ distance between embeddings, the verification process instead focuses on finding the maximum verifiable radius under the $\ell_{\infty}$ norm, since formal verification over the $\ell_{2}$ domain is not directly possible using zonotopes. The authors intentionally adopt the $\ell_{2}$ distance for pre-training because it minimizes the overall distance between points, rather than just the maximum per-dimension difference. This encourages the model to learn more generalized and robust representations. Furthermore, the authors observed in their experiments that after pre-training, the $\ell_{2}$ distance between semantically similar words (such as gendered word pairs) is often slightly larger than their corresponding $\ell_{\infty}$ distance (as can be seen in Section C.3). This supports the idea that utilizing the $\ell_{2}$ norm benefits downstream verification.
@@ B.3 Embedding model pre-training data construction
Given these limitations and the impracticality of verifying large-scale contextualized models like BERT (Shi et al., 2020; Bonaert et al., 2021), we pre-train a Keras embedding layer (Ketkar and Ketkar, 2017). To construct an effective training set for contrastive loss optimization, we curated word pairs based on the semantic gender relationship using a heuristic approach. Each male-labeled word was explicitly paired with the terms woman and female to anchor the gender axis, and further matched with a diverse sample of 50-100 female-labeled words to form similar (label=0) pairs. This ensures that embeddings of gender-related terms are pulled closer in the vector space. To balance the contrastive objective, we generated dissimilar (label=1) pairs by selecting words from a broad English vocabulary that have no direct lexical or semantic relationship (e.g., synonyms, derivations) to any gender-specific term. This distinction was enforced by filtering out related terms using WordNet relations such as synonyms, hypernyms, and derivationally related forms.
@@
To enhance the general semantic capacity of the embedding space, particularly because the embedding layer is later frozen during task-specific backpropagation, we conducted an auxiliary contrastive training phase focused on common semantic relationships such as synonymy (This general semantic training step proved essential, as model performance degraded significantly in its absence. This outcome is expected, given that the embedding layer, when integrated within the transformer-based architecture, is frozen during downstream training and therefore does not receive gradient updates to refine its semantic representations). Using WordNet, we generated word pairs labeled as similar (label=0) by extracting synonym pairs, while carefully excluding any gender-specific terms to preserve the integrity of gender representations learned in earlier pre-training stages, thereby maintaining the intended semantic distances between gendered terms. To form the dissimilar (label=1) pairs, we randomly sampled words from the general vocabulary that were not semantically related (each non-gendered word is paired with all valid WordNet synonyms, excluding any gendered terms, resulting in a variable number of synonym pairs per word. To ensure a balanced dataset, an equal number of dissimilar word pairs is generated, maintaining a one-to-one ratio with the total number of synonym pairs), again ensuring that no gender-specific tokens were involved. This allowed the model to learn to cluster semantically similar words and push apart unrelated ones. By training the embedding space on these general lexical relationships, we improve its ability to capture broader semantic structure, which is especially important since the embeddings are utilized as fixed features in downstream transformer-based language model text classification tasks. This step ensures that the embeddings are not only gender-aware but also semantically meaningful in a general linguistic context.
@@
For pre-training the toxicity-related embedding model, a similar pre-training strategy was employed—using a contrastive learning framework to pull semantically similar words closer and push dissimilar words apart (to handle rare or profane words not explicitly found in the tokenizer vocabulary, we use BERT’s WordPiece-based subword tokenization. This strategy breaks out-of-vocabulary terms into smaller known units (e.g., mother-fucker becomes [mother, -, fuck, ##er]), allowing consistent representation throughout the embedding layer pre-training and language model training phases without full word-level coverage. These subwords are embedded and mean-pooled, enabling the model to capture semantic information even for obfuscated or stylized expressions). The strength of the pulling and pushing mechanisms (i.e., the values of $\alpha$ and the margin), along with hyperparameters, can be found in Section C.1. The distance between gendered words, toxicity words, synonyms, and unrelated words before and after pre-training of embedding can be found in Section C.3.
@@ B.4 Adult census dataset preprocessing
Transformers require one-dimensional token sequences as input, whereas tabular data is inherently two-dimensional (Badaro et al., 2023); however, the demographic information in the Adult Census dataset is structured in a tabular format, with each attribute represented as a separate column. To bridge this gap, the data and its contextual structure must be serialized before being processed by the model (Badaro et al., 2023). A critical step in adapting tabular data for transformer-based models is converting the tabular structure into a linearized, one-dimensional format. Common strategies include: column-wise concatenation of values, insertion of special separator tokens between fields (Badaro et al., 2023; Suhara et al., 2022), and use of text templates to represent each row as a structured sentence (Chen et al., 2020d; Suadaa et al., 2021; Chen et al., 2020c). More recent approaches move beyond predefined templates by generating natural language sentences directly from tabular data. This is achieved by fine-tuning sequence-to-sequence models such as T5 (Raffel et al., 2020) or GPT (Radford et al., 2019), as demonstrated in prior works (Thorne et al., 2021; Neeraja et al., 2021). We adopt simple concatenation for its straightforward implementation and effectiveness in our setting.
To adapt the Adult Census dataset for use with a sentence-to-label transformer-based binary classification model, we converted the structured tabular data into natural language sentences. The original dataset contains both numerical and categorical features, along with information indicating whether an individual’s income exceeds \$50K. Since transformer models (e.g., BERT) operate on text sequences, we constructed a heuristic template that combines relevant features into coherent sentences. Specifically, attributes such as workclass, education, occupation, sex, marital status, relationship, race, native country, and hours per week were embedded into a natural language sentence format. The salary attribute, serving as the ground truth label for prediction, was intentionally excluded from the sentence construction. This transformation preserves the semantic content of the original features while enabling the use of sentence-based binary text classification models on structured tabular data. An example tabular data instance before and after sentence construction can be found in Section B.4 (the feature ”fnlwgt”, along with several other complex numerical attributes, was excluded from this experiment, as the authors deliberately chose to focus on leveraging textual, simple numerical, and categorical features for model training and representation).
@@ B.5 Data augmentation
Text classification is a fundamental task in natural language processing, and achieving high performance often depends heavily on the size and quality of the training data (Wei and Zou, 2019; Kobayashi, 2018). Automatic data augmentation is widely employed in machine learning and computer vision, among other domains, to enhance model robustness, particularly when working with limited datasets (Simard et al., 1998; Szegedy et al., 2015; Krizhevsky et al., 2012; Cui et al., 2015; Ko et al., 2015).
A commonly used data augmentation strategy is synonym replacement, where selected words are substituted with their synonyms. This can be achieved using handcrafted lexical resources like WordNet (Miller, 1995; Zhang et al., 2015), word similarity computations (Wang and Yang, 2015), or contextual embedding-based methods (Kobayashi, 2018).
We employed a simple WordNet-based synonym replacement strategy to augment the dataset in the salary prediction experiments, aiming to increase data volume and enhance classification accuracy. This approach led to a notable improvement in test accuracy, rising from 80% to around 90% (we utilized straightforward WordNet synonym replacement, as empirical observations indicated that this method is already effective enough in improving model performance).
@@ C.1 Evaluation details
The selection of margin and $\alpha$ values is based on empirical observation. The goal is not only to ensure that gendered words are embedded in close proximity but also to promote the clustering of synonyms while pushing antonyms farther apart, without compromising the relative closeness of gendered word pairs. The margin values are also selected empirically based on the nuanced relationship between the embedding space and the resulting verifiable radius. A detailed study of this relationship is provided in Section C.3. The loss function used in this study is focal loss (Lin et al., 2017), a training-time calibration technique that directly integrates class imbalance handling into the learning process. It works by adapting the loss function to emphasize challenging, misclassified samples.
To evaluate the fairness score of the models, we use the MATLAB toolbox CORA (Althoff, 2015), which enables formal verification using set-based computing (Kochdumper et al., 2023; Ladner and Althoff, 2023; Bonaert et al., 2021). Please visit Appendix A for a detailed explanation of this process using zonotopes.
@@
[…]
Model $\alpha$ Margin $\alpha$ (Gen. Sem. Train.) Margin (Gen. Sem. Train.) Epoch Epoch (Gen. Sem. Train.)
@@
[…]
Model $\alpha$ Margin $\alpha$ (Gen. Sem. Train.) Margin (Gen. Sem. Train.) Epoch Epoch (Gen. Sem. Train.)
@@
[…]
Model Vocab (BERT) Seq. Len. Embed Dim Heads FF Dim Dropout Epochs Batch
@@
[…]
Model Vocab Seq. Len. Embed Dim Heads FF Dim Dropout Epochs Batch
@@ C.2 Sentence-wise maximum verifiable perturbation radius
[…]
@@
Sentences 1–25 Sentences 26–50
Sentence Max verifiable noise radius ($\uparrow$) Sentence Max verifiable noise radius ($\uparrow$)
Noise Radii Sentence 1 8.6079e-04 Sentence 26 8.2806e-04
Sentence 2 4.9749e-04 Sentence 27 7.7405e-04
Sentence 3 2.9457e-04 Sentence 28 8.5915e-04
Sentence 4 8.8533e-04 Sentence 29 8.5915e-04
Sentence 5 6.5950e-04 Sentence 30 8.7715e-04
Sentence 6 8.5915e-04 Sentence 31 8.5915e-04
Sentence 7 6.5459e-04 Sentence 32 3.1748e-04
Sentence 8 6.8896e-04 Sentence 33 9.3279e-04
Sentence 9 7.2169e-04 Sentence 34 6.8896e-04
Sentence 10 2.8147e-04 Sentence 35 5.9895e-04
Sentence 11 6.8896e-04 Sentence 36 7.8060e-04
Sentence 12 8.5915e-04 Sentence 37 7.2823e-04
Sentence 13 8.5915e-04 Sentence 38 2.9457e-04
Sentence 14 8.5751e-04 Sentence 39 3.5512e-04
Sentence 15 6.8896e-04 Sentence 40 8.2806e-04
Sentence 16 9.4752e-04 Sentence 41 8.2806e-04
Sentence 17 4.7949e-04 Sentence 42 8.1987e-04
Sentence 18 8.2806e-04 Sentence 43 8.2806e-04
Sentence 19 5.8586e-04 Sentence 44 7.9696e-04
Sentence 20 8.0515e-04 Sentence 45 1.9638e-04
Sentence 21 2.9457e-04 Sentence 46 9.1152e-04
Sentence 22 6.1368e-04 Sentence 47 8.3624e-04
Sentence 23 6.7586e-04 Sentence 48 8.5915e-04
Sentence 24 8.5915e-04 Sentence 49 8.2806e-04
Sentence 25 6.9878e-04 Sentence 50 7.6751e-04
@@
[…]
Embedding Bounds Minimum ($\inf$) -0.838544
Maximum ($\sup$) 0.865149
Range ($\sup-\inf$) 1.703692
Sentences 1–25 Sentences 26–50
Sentence Max verifiable noise radius ($\uparrow$) Sentence Max verifiable noise radius ($\uparrow$)
Noise Radii Sentence 1 4.2090e-03 Sentence 26 4.3820e-03
Sentence 2 6.3423e-03 Sentence 27 4.9913e-03
Sentence 3 4.6858e-03 Sentence 28 6.2285e-03
Sentence 4 4.0839e-03 Sentence 29 4.0562e-03
Sentence 5 4.4446e-03 Sentence 30 3.5509e-03
Sentence 6 4.8865e-03 Sentence 31 4.3999e-03
Sentence 7 4.8061e-03 Sentence 32 3.3641e-03
Sentence 8 5.5104e-03 Sentence 33 3.0164e-03
Sentence 9 3.8215e-03 Sentence 34 3.0164e-03
Sentence 10 4.6940e-03 Sentence 35 4.5656e-03
Sentence 11 3.7451e-04 Sentence 36 3.3454e-03
Sentence 12 5.2959e-03 Sentence 37 5.4438e-03
Sentence 13 4.6745e-03 Sentence 38 4.0497e-03
Sentence 14 1.4160e-03 Sentence 39 5.1343e-03
Sentence 15 5.0148e-03 Sentence 40 3.3015e-03
Sentence 16 4.7354e-03 Sentence 41 4.4242e-03
Sentence 17 5.2748e-03 Sentence 42 3.9230e-03
Sentence 18 4.7403e-03 Sentence 43 4.0457e-03
Sentence 19 5.7736e-03 Sentence 44 7.1896e-04
Sentence 20 2.9083e-03 Sentence 45 4.3341e-03
Sentence 21 2.8401e-03 Sentence 46 3.1326e-03
Sentence 22 5.3471e-03 Sentence 47 4.9742e-03
Sentence 23 5.0067e-03 Sentence 48 4.8288e-03
Sentence 24 4.7801e-03 Sentence 49 4.5640e-03
Sentence 25 3.5371e-03 Sentence 50 5.7557e-03